40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m elog  Not logged in ELOG logo
Entry  Tue Sep 22 23:16:54 2020, Koji, Update, General, HEPA Inspection 11x
    Reply  Tue Sep 22 23:17:51 2020, Koji, Update, General, HEPA Inspection 6x
       Reply  Wed Sep 30 17:12:24 2020, gautam, Update, General, HEPA blower startup capacitor replacement 
          Reply  Wed Sep 30 19:45:56 2020, rana, Update, General, HEPA blower startup capacitor replacement 
Message ID: 15597     Entry time: Tue Sep 22 23:16:54 2020     Reply to this: 15598
Author: Koji 
Type: Update 
Category: General 
Subject: HEPA Inspection 

Gautam reported that the PSL HEPA stopped running (ELOG 15592). So I came in today and started troubleshooting.

It looks like that the AC power reaches the motors. However, both motors do not run. It looks like the problem exists in the capacitors, the motors, or both.
Parts specs can be found in the next ELOG.


Attachment 1 is the connection diagram of the HEPA. The AC power is distributed by the breaker panel. The PSL HEPA is assigned to use M22 breaker (Attachment 2). I checked the breaker switch and it was (and is) ON. The power goes to the junction box above the enclosure (Attachment 3). A couple of wires goes to the HEPA switch (right above the enclosure light switch) and the output goes to the variac. The inside of the junction box looked like this (Attachment 4).

By the way, the wires were just twisted and screwed into a metal threaded (but isolated) caps (Attachment 5). Is this legit? Shouldn't we use stronger crimping? Anyway, there was nothing wrong with the caps w.r.t the connection for now.

I could easily trace the power up to the variac. The variac output was just fine (Attachment 6). The cord goes from the variac to the junction box (and then HEPAs) looked scorched. The connection from the plug to HEPAs was still OK, but this should be eventually replaced. Right now the cable was unplugged after the following tests for the safety reason.

The junction box for each HEPA unit was opened to check the voltage. The supply voltage came to the junction boxes and it was just fine. In Attachments 8 & 9, the voltages look low but this is because I just turned the variac only a little.

At the (main) junction box, the resistances of the HEPAs were checked with the Fluke. As the HEPA units are connected to the AC in parallel, the resistances were individually checked as follows.

South HEPA SW North HEPA SW Resistance
OFF OFF High
OFF LO 5 Ohm
OFF HIGH 7 Ohm
LO OFF 7 Ohm
HIGH OFF 5 Ohm

The coils were not disconnected (... I wonder if the wiring of South HEPA was flipped? But this is not the main issue right now.)
 

By removing the pre-filters, the motors were inspected Attachments 10 & 11. At least the north HEPA motor was warm, indicating there was some current before. A capacitor was connected per motor. When the variac was tuned up a bit, one side of the capacitor could see the voltage. I could not judge which has the issue between the capacitor and the motor.

Attachment 1: 0_PSL_HEPA.pdf  22 kB  Uploaded Wed Sep 23 00:19:09 2020  | Show | Hide all | Show all
Attachment 2: 1_Breaker_Panel.JPG  596 kB  Uploaded Wed Sep 23 00:19:21 2020  | Hide | Hide all | Show all
1_Breaker_Panel.JPG
Attachment 3: 2_Junction_Box.JPG  775 kB  Uploaded Wed Sep 23 00:19:21 2020  | Hide | Hide all | Show all
2_Junction_Box.JPG
Attachment 4: 3_Junction_Box_Inside.JPG  4.160 MB  Uploaded Wed Sep 23 00:19:21 2020  | Show | Hide all | Show all
Attachment 5: 4_Junction_Box_Inside_2.JPG  3.869 MB  Uploaded Wed Sep 23 00:19:21 2020  | Hide | Hide all | Show all
4_Junction_Box_Inside_2.JPG
Attachment 6: 5_Variac_100%.JPG  4.483 MB  Uploaded Wed Sep 23 00:19:21 2020  | Hide | Hide all | Show all
5_Variac_100%.JPG
Attachment 7: 6_VariAC_to_HEPA.JPG  470 kB  Uploaded Wed Sep 23 00:19:21 2020  | Hide | Hide all | Show all
6_VariAC_to_HEPA.JPG
Attachment 8: 7_North_HEPA_IN.JPG  1.188 MB  Uploaded Wed Sep 23 00:19:21 2020  | Show | Hide all | Show all
Attachment 9: 8_South_HEPA_IN.JPG  823 kB  Uploaded Wed Sep 23 00:19:21 2020  | Hide | Hide all | Show all
8_South_HEPA_IN.JPG
Attachment 10: 9_North_Prefilter_Removed.JPG  4.283 MB  Uploaded Wed Sep 23 00:19:21 2020  | Show | Hide all | Show all
Attachment 11: 10_South_Prefilter_Removed.JPG  4.088 MB  Uploaded Wed Sep 23 00:19:21 2020  | Hide | Hide all | Show all
10_South_Prefilter_Removed.JPG
ELOG V3.1.3-